# Data Publishing in Mechanics and Dynamics: Challenges, Guidelines, and Examples from Engineering Design

### v1

@@ 6 Conclusions and Outlook
Data science, machine learning, and artificial intelligence have in recent years expanded from first success stories in computer vision and natural language processing to progress in natural sciences and engineering. Regarding the latter, artificial intelligence also starts to have a lasting impact on engineering design, and will transform how humanity conceives new machines, vehicles, robots, etc. As this article has outlined, engineering design brings about characteristic perspectives and challenges of which not every aspect and challenge are present for the currently most successful applications of machine learning. In particular, this includes the challenge of evaluation, as for design problems, there usually is no known best or true solution relative to which the quality of results can be gauged. Also, computationally evaluating the quality of a proposed design can be very expensive, e.g., by performing expert-driven and computationally highly demanding simulations, whereas data from experiments is usually not yet available in early design stages or very scarce. As a consequence, the generation of large datasets becomes prohibitively expensive. Even worse, these simulations often benefit from high-performance computing architectures different than the ones most fitting for, e.g., deep learning as engineering simulations typically rely more significantly on general-purpose processing. In the future, this may be alleviated, e.g., through simulators that, themselves, are based on machine-learning approaches. Nevertheless, when proposing and evaluating radically new designs, there remains the risk that such machine-learned simulators would be evaluated in regions where they have no prior knowledge, which may yield arbitrarily unreliable results, or require currently unseen generalization. Still, as the examples provided in this paper show, today’s methods and state of the art already allow to publish and use data and data-based approaches in the field of engineering design, giving first answers of how to overcome associated challenges. However, clearly, researchers in engineering are only at the beginning of transforming engineering design toward a data-driven future. And one of the most interesting aspects will be whether this will lead to a future full of auto-designed, highly task-specific machinery that can be extremely efficient and accurate due to designs perfected for the specific task, not limited by the amount of time, money, and knowledge available in the form of human design engineers.
@@
However, as publication habits change to include ever more data and code, also the question arises whether reviewing processes need to adapt. Does the reviewer of the future need to rate the interpretability of supplied data and ensure that future, competitive approaches can be readily and fairly evaluated? Many more questions of this kind will accompany the future of publication habits in mechanics, dynamics, and engineering design.
Acknowledgements
@@
All authors acknowledge the support by the German Research Foundation (Deutsche Forschungsgemeinschaft, DFG) within the Priority Program (DFG SPP) 2353 ”Daring More Intelligence: Design Assistants in Mechanics and Dynamics”. P. Eberhard acknowledges support under DFG Grant No. 501890093 ”Daring More Intelligence – Design Assistants in Mechanics and Dynamics (SPP 2353) - Coordination Proposal”, and H. Ebel is grateful for the past support from the same grant. P. Eberhard and B. Röder acknowledge support under DFG Grant No. 501840485. J. van Delden and T. Lüddecke acknowledge the support by the Deutsche Forschungsgemeinschaft (German Research Foundation), project number 501927736, within the DFG Priority Program 2353. They also acknowledge the computing time made available to them on the high-performance computers HLRN-IV at GWDG at the NHR Centers NHR@Göttingen. These centers are jointly supported by the German Federal Ministry of Education and Research and the German state governments participating in the NHR (www.nhr-verein.de/unsere-partner). A. Borse, R. Gulakala, and M. Stoffel gratefully acknowledge the financial support provided by Deutsche Forschungsgemeinschaft Priority Program SPP 2353 (DFG Grant No. STO 469/16-1). M. Stender and M. Yadav acknowledge the support by the Deutsche Forschungsgemeinschaft under grant 501847579. L. Schindler and K. M. de Payrebrune gratefully acknowledge the funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 501861263 – SPP2353. M. Raff and C. D. Remy acknowledge the funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 501862165. M. Raff further acknowledges the support through the International Max Planck Research School for Intelligent Systems (IMPRS-IS).

### v2

@@ 6 Conclusions and Outlook
Data science, machine learning, and artificial intelligence have in recent years expanded from first success stories in computer vision and natural language processing to progress in natural sciences and engineering. Regarding the latter, artificial intelligence also starts to have a lasting impact on engineering design, and will transform how humanity conceives new machines, vehicles, robots, etc. As this article has outlined, engineering design brings about characteristic perspectives and challenges of which not every aspect and challenge are present for the currently most successful applications of machine learning. In particular, this includes the challenge of evaluation, as for design problems, there usually is no known best or true solution relative to which the quality of results can be gauged. Also, computationally evaluating the quality of a proposed design can be very expensive, e.g., by performing expert-driven and computationally highly demanding simulations, whereas data from experiments is usually not yet available in early design stages or very scarce. As a consequence, the generation of large datasets becomes prohibitively expensive. Even worse, these simulations often benefit from high-performance computing architectures different than the ones most fitting for, e.g., deep learning as engineering simulations typically rely more significantly on general-purpose processing. In the future, this may be alleviated, e.g., through simulators that, themselves, are based on machine-learning approaches. Nevertheless, when proposing and evaluating radically new designs, there remains the risk that such machine-learned simulators would be evaluated in regions where they have no prior knowledge, which may yield arbitrarily unreliable results, or require currently unseen generalization. Still, as the examples provided in this paper show, today’s methods and state of the art already allow to publish and use data and data-based approaches in the field of engineering design, giving first answers of how to overcome associated challenges. Clearly, researchers in engineering are only at the beginning of transforming engineering design toward a data-driven future. And one of the most interesting aspects will be whether this will lead to a future full of auto-designed, highly task-specific machinery that can be extremely efficient and accurate due to designs perfected for the specific task, not limited by the amount of time, money, and knowledge available in the form of human design engineers.
@@
However, as publication habits change to include ever more data and code, also the question arises whether reviewing processes need to adapt. Does the reviewer of the future need to rate the interpretability of supplied data and ensure that future, competitive approaches can be readily and fairly evaluated? Will there be scientifically peer-reviewed publications of data, without accompanying paper and with only the necessary data usage information and processing code, for which citations will count toward key performance indicators for scientific careers? Many more questions of this kind will accompany the future of publication habits in mechanics, dynamics, and engineering design.
Acknowledgments
@@
All authors acknowledge the support by the German Research Foundation (Deutsche Forschungsgemeinschaft, DFG) within the Priority Program (DFG SPP) 2353 ”Daring More Intelligence: Design Assistants in Mechanics and Dynamics”. P. Eberhard acknowledges support under DFG Grant No. 501890093 ”Daring More Intelligence – Design Assistants in Mechanics and Dynamics (SPP 2353) - Coordination Proposal”, and H. Ebel is grateful for the past support from the same grant. P. Eberhard and B. Röder acknowledge support under DFG Grant No. 501840485. J. van Delden and T. Lüddecke acknowledge the support by the Deutsche Forschungsgemeinschaft (German Research Foundation), project number 501927736, within the DFG Priority Program 2353. They also acknowledge the computing time made available to them on the high-performance computers HLRN-IV at GWDG at the NHR Centers NHR@Göttingen. These centers are jointly supported by the German Federal Ministry of Education and Research and the German state governments participating in the NHR (www.nhr-verein.de/unsere-partner). A. Borse, R. Gulakala, and M. Stoffel gratefully acknowledge the financial support provided by Deutsche Forschungsgemeinschaft Priority Program SPP 2353 (DFG Grant No. STO 469/16-1). M. Stender and M. Yadav acknowledge the support by the Deutsche Forschungsgemeinschaft under grant 501847579. L. Schindler and K. M. de Payrebrune gratefully acknowledge the funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant number 501861263. M. Raff and C. D. Remy acknowledge the funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 501862165. M. Raff further acknowledges the support through the International Max Planck Research School for Intelligent Systems (IMPRS-IS). The contribution of R. Raj, T. Rentschler, A. Tismer, and S. Riedelbauch was funded by the German Research Foundation (DFG) (Grant/Award Number: 501932169) and the corresponding simulations were performed on the national supercomputer HPE Apollo Hawk cluster at the High-Performance Computing Center Stuttgart (HLRS) under the grant number parDtOOMeetsCanada/44196. The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.